OpenStudy (anonymous):

Solve the following equation for x: -4 sqrt(x+9) = 20

OpenStudy (anonymous):

i think this might be an extraneous solution

OpenStudy (mind2000):

SQRT(X-8)=X-10 SQUARE BOTH SIDES X-8=X^2-20X+100 X^2-20X-X+100+8=0 X^2-21X+108=0 (X-12)(X-9)=0 X-12=0 X=12 ANSWER X-9=0 X=9 ANSWER.

OpenStudy (anonymous):